OpenStudy (anonymous):

Calculate the equation of the line between the points (-6, -1) and (2, -1)

OpenStudy (kainui):

What is the standard equation of a line? y=mx+b. m is the slope, it is rise/run and b is the y-intercept. Try drawing it out to see if this helps.

OpenStudy (anonymous):

I tried yet I dont understand.

OpenStudy (kainui):

Well, what does your picture look like? How much do you rise and how much do you run to get from one point to the next? Where on the y-axis is the line touching at x=0?

OpenStudy (anonymous):

Some what above -1?

OpenStudy (anonymous):

(-6, -1) and (2, -1) First find the slope of the line: y - y over x - x Once you find the slope try putting it into point-slope form: y - y1 = m(x - x1)

OpenStudy (anonymous):

After that you can change it easily from point-slope to standard form.

OpenStudy (anonymous):

Im completely lost.

OpenStudy (anonymous):

Do you want to do it the way I mentioned above? Step by step?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Ok, so first lets find the slope: The formula to find slope is the difference between the y points over the difference between the x points.

OpenStudy (anonymous):

So the one furthest to the right is Y2? And X2?

OpenStudy (anonymous):

Look at your given points: (-6, -1) and (2, -1) y values: -1 and -1 x values: 2 and -6

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So find the difference between the y-values: -1 - (-1)

OpenStudy (kainui):

Jeremiah, I know this might take 30 minutes, but really play around with a graph of y based on x being multiplied with numbers and being added to numbers. Then you'll understand why y=mx+b. What is the difference between the graphs y=2x and y=x? What is the difference between y=x+1 and y=x? If you spend the time to figure this out you will understand and maybe even be able to answer your own questions! =D

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I agree with @Kainui when you put these points on a graph it is definately easier to spot out the equation:) Check again: Try -1 + 1 (two - signs equal a +)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

sorry i did that and I thoughh it said 1+1

OpenStudy (anonymous):

Yes. And 2 + 6?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

0/8

OpenStudy (anonymous):

Exactly! You see how there is a 0 in the numerator. Remember rise/run: so the line doesn't rise. The slope is 0.

OpenStudy (anonymous):

B is slope right or is it m?

OpenStudy (anonymous):

The line is a horizontal line. m = slope b = y-intercept

OpenStudy (anonymous):

So whats y and x?

OpenStudy (anonymous):

Ok so now we have the slope and we can make the point-slope equation, This is quick and easy:)

OpenStudy (anonymous):

Pick one of the two original points (-6, -1) or (2, -1)

OpenStudy (anonymous):

Whats the Y intercept?

OpenStudy (anonymous):

y-intercept is where the line crosses the y-axis. But in order to get to standard form (y = mx + b) you should put an equation is point-slope form(y - y1 = m(x - x1) So you pick one point. Lets go with (-6, -1). Now just substitute the values into the equation: (-6, -1) m= slope = 0 y - (-1) = 0(x - (-6)) y + 1 = 0(x + 6)

OpenStudy (anonymous):

**The y1 and x1 are where you put in the values of x and y

OpenStudy (anonymous):

so the final equation Im geussing would end up bein y=-8x-1?

OpenStudy (anonymous):

no. Remember that the m in y = mx + b is the slope. And we found that the slope was 0. But your y-inercept(b) is correct:) Try simplifying this: First you must distribute the 0 into the x and the 6... y + 1 = 0(x + 6) y + 1 = (0 * x) + (0 * 6) Keep in mind that you can never elliminate the x; 0 * x would be 0x

OpenStudy (anonymous):

so what would it be?

OpenStudy (anonymous):

So it would be: y + 1 = 0x Now you just want to leave the y alone on one side so subtract the 1 from both sides of the equation: y + 1 - 1 = 0x -1 y = 0x - 1 Final answer: The equation of the line between the points (-6, -1) and (2, -1) would be: \[y = 0x + 1\]

OpenStudy (anonymous):

Thats not one of the choices though

OpenStudy (anonymous):

What are your choices? Sorry it took so long to explain:) It looks complicated at first but it gets a lot easier after you do it a lot.

OpenStudy (anonymous):

a. y = -1 b. x = -1 c. y = -8x - 1 d. y = -x - 8

OpenStudy (anonymous):

A) y = -1 Sorry I had forgotten that when a line is horizontal the answer is just y = and the y-intercept.... Whenever you want to check an equation i suggest you use this: https://www.desmos.com/calculator Just put in the equation(doesn't matter the format) and it shows the line. Out of curiosity, are you in FLVS?

OpenStudy (anonymous):

Oh no Im in BYU

OpenStudy (anonymous):

Oh ok:) I was asking cuz I took Algebra in FLVS last semester; and this was pretty much what I came to ask at openstudy every day until I got the hang of it:D

OpenStudy (anonymous):

Hope I helped, as confusing as it might have been...:) And I hope you do well on your classes!